# A mixed logic with binary operators

### v1

@@ Abstract.
In previous work (Düntsch et al., 2023) we introduced and examined the class of betweenness algebras. In the current paper we study a larger class of algebras with binary operators of possibility and sufficiency, the weak mixed algebras. Furthermore, we develop a system of logic with two binary modalities, sound and complete with respect to the class of frames closely related to the aforementioned algebras, and we prove an embedding theorem which solves an open problem from (Düntsch et al., 2023).
@@ 1. Introduction
Possibility-sufficiency algebras (PS-algebras), as Boolean algebras expanded with unary operators of possibility and sufficiency, were first studied in (Düntsch and Orłowska, 1999). The latter operator is an algebraic counterpart of the window modality, whose provenance can be traced back to the works of Humberstone (1983), Gargov et al. (1987), Goranko (1988) and also Van Benthem (1979) in the framework of deontic logic. The theory of PS-algebras was refined and developed further by Düntsch and Orłowska (2001) and Düntsch et al. (2017), with particular emphasis on two subclasses of PS-algebras: mixed and weak mixed algebras, which force a relationship between the—otherwise independent—operators of possibility and sufficiency.
@@
Building on the aforementioned works, we initiated in (Düntsch et al., 2023) a theory of betweenness algebras (b-algebras), which were thought of as an approach to study most general algebraic and logical properties of the ternary relation of betweenness well-known from ordered geometry. Being ternary, betweenness calls for binary operators, and for this reason, our PS-algebras are Boolean algebras with binary operators of possibility and sufficiency. To guarantee the interplay between the two operators, we focused on weak mixed algebras, which we enriched with equalities expressing the basic and well-motivated properties of betweenness. The approach turned out to be fruitful and allowed for the development of both representation and canonical extensions of algebras arising from betweenness relations. For these, the canonical frames of b-algebras are triples of the form $\langle W,R,S\rangle$, where $S$ and $R$ are ternary relations on $W$ obtained in the standard way from sufficiency and possibility operators, respectively, with $S\subseteq R$. In a certain sense, $R$ and $S$ together simulate the betweenness relation. We showed that this cannot be avoided, in the sense that, in general, the canonical frame of a b-algebra cannot have just one ternary relation which is a betweenness. This was a negative result, as the starting point of our research were precisely frames $\langle W,B\rangle$, where $B$ is a betweenness, and the properties of the complex algebras of such frames gave rise to the axiomatization presented in the paper. However, the negative finding did not exclude existence of a more general ternary relation that could allow for a representation theorem for b-algebras, and finding such a relation was formulated as problem (1) in (Düntsch et al., 2023).
@@
With the problem in focus, we concentrate below on a general study of frames $\langle W,R,S\rangle$ with $S\subseteq R$ (wMIA frames). We develop a logic $K^{\#}$ with binary operators of possibility and sufficiency (window modality) that is sound and complete with respect to the class of wMIA frames. To this end, we adapt for our needs the copying method from (Vakarelov, 1989, Gargov et al., 1987), which allows for the transformation of wMIA frames into frames with a single ternary relation. Further, we study algebraic models of $K^{\#}$, i.e., PS-algebras determined by an equational theory $\Sigma$. We prove that the variety $\operatorname{\mathbf{V}}(\Sigma)$ is precisely the variety generated by the class of weak mixed algebras (which does not form a variety itself).
@@ 2.1. Boolean algebras
[…]
@@ 2.2. PS-algebras
We prepare the description of the congruences on a PS-algebra $\mathfrak{A}$ with a few known facts. The Boolean congruences on $A$ were described earlier in Lemma 2.1. Congruences with respect to a binary necessity operator were characterized by Venema (2007).
Lemma 2.2 (Venema, 2007, Theorem 29).
@@
To apply these results, we first define for a sufficiency operator $g:A\times A\to A$ a mapping $g^{*}:A\times A\to A$ by $g^{*}(a,b)=g(-a,-b)$. The operator $g^{*}$ is a binary necessity operator, and a Boolean congruence preserves $g$ if and only if it preserves $g^{*}$, since $g$ and $g^{*}$ are mutually term definable. We now define an auxiliary mapping $u\colon A\times A\to A$ [^1] The name refers to the universal modality which is related to the unary discriminator, see e.g. Gargov et al. (1987), Goranko and Passy (1992).
@@ 2.3. Complex algebras and canonical frames
The canonical frame of $\mathfrak{A}$ is the structure $\operatorname{{\mathsf{Cf}}}(\mathfrak{A})\coloneqq\langle\operatorname{Ult}(A),R_{f},S_{g}\rangle$, where $R_{f},S_{g}$ are ternary relations on $\operatorname{Ult}(A)$ defined by
@@
($\supseteq$) Let $\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}\in\operatorname{Ult}(A)$, and $a\in\mathscr{U}_{1},b\in\mathscr{U}_{3}$. If $R_{f}(\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3})$, then, $f(a,b)\in\mathscr{U}_{2}$, and it follows that $h(a,b)=f(a,b)+-g(a,b)\in\mathscr{U}_{2}$. If $\langle\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}\rangle$ is not an element of $S_{g}$, then $g[\mathscr{U}_{1}\times\mathscr{U}_{3}]\cap\mathscr{U}_{2}=\emptyset$, hence, $g(a,b)\not\in\mathscr{U}_{2}$ which implies $-g(a,b)\in\mathscr{U}_{2}$. It follows that $h(a,b)=f(a,b)+-g(a,b)\in\mathscr{U}_{2}$. ∎
@@ 3. The binary logic $K^{\#}$
$[U]](\varphi,\psi)\coloneqq\Box(\varphi,\psi)\wedge\boxbar(\neg\varphi,\neg\psi)\,.$
@@
$K^{\#}$ is the smallest set of formulas that contains propositional tautologies, all instances of (M1)–(M4), ($\mathrm{C}_{U}$)–($\mathrm{B}_{U}$) and is closed under modus ponens and uniform substitution.
@@ 3.1. Relational semantics
(wMIA) $(\forall a,b\in A)\,[a\neq\mathbf{0}\text{ and }b\neq\mathbf{0}\Rightarrow g(a,b)\leq f(a,b)]\,.$
@@
Notation not explained here can be found in (Blackburn et al., 2001, Section 5), Düntsch et al. (2023), and Burris and Sankappanavar (2012), in particular, II §10.
@@
[…]
@@ 3.2. The canonical frame for $\boldsymbol{K^{\#}}$
$[U]](\langle U\rangle\!\rangle(\langle U\rangle\!\rangle(\psi,\top),\top),\bot)\in y$
@@
$[U]](\langle U\rangle\!\rangle(\psi,\top),\bot)\in y\,.$
By the same argument $[U]](\langle U\rangle\!\rangle(\varphi,\top),\bot)\in y$. Since
$[U]](\langle U\rangle\!\rangle(\varphi,\top),\bot)\land[U]](\langle U\rangle\!\rangle(\psi,\top),\bot)\Leftrightarrow[U]](\langle U\rangle\!\rangle(\varphi,\top)\land\langle U\rangle\!\rangle(\psi,\top),\bot),$
@@
$[U]](\langle U\rangle\!\rangle(\varphi,\psi),\bot)\in y,\quad\text{i.e.,}\quad\neg\langle U\rangle\!\rangle(\neg\langle U\rangle\!\rangle(\varphi,\psi),\top)\in y\,.$
@@ 3.3. Special frames and models
Our next aim is to construct—given $\langle W,R,S,v\rangle$—a special model $\langle\underline{W},\underline{R},\underline{S},\underline{v}\rangle$, that is modally equivalent to $\langle W,R,S,v\rangle$. The idea of the construction of $\underline{R}$ and $\underline{S}$ is to “separate” triples $\langle x,y,z\rangle$ which are in the intersection of $R$ and $-S$, i.e. to remove those triples which prevent $R=S$ (see Figure 1).
@@
[…]
@@
()superset-of-or-equals($\supseteq$)( ⊇ ) If $\tau\in m(W^{3}\setminus P)$, then there is $\langle x,y,z\rangle\notin P$ such that $\tau\in c(x,y,z)$. If there were $\langle r,s,t\rangle\in m(P)$ such that $\tau\in c(j(r),j(s),j(t))$, then $P$ and $W^{3}\setminus P$ would have a non-empty intersection, which is impossible.
@@
The mixture $m(\mathfrak{F})$ of any wMIA frame $\mathfrak{F}$ can be turned into a special frame via the following construction of $\underline{R}$ and $-\underline{S}$ (below, $-R$ and $-S$ are the complements of the two relations in $W^{3}$):
@@
Figure 2. $m(R)\setminus m(S)$ is equal to the union of $[m(R)\setminus m(S)]\setminus(R\cup R^{\prime})$ and $(R\cup R^{\prime})\setminus(S\cup S^{\prime})$.
@@
$[(v(\varphi)\cup v^{\prime}(\varphi))\times\{s\}\times(v(\psi)\cup v^{\prime}(\psi))]\cap\underline{R}\neq\emptyset\,.$
@@ 4. The class $\operatorname{\mathbf{V}}(\mathbf{wMIA})$
In the case of unary modalities investigated in (Düntsch et al., 2017) a unary PS-algebra $\langle A,f,g\rangle$ is a weak MIA if and only if the mapping defined by $u^{\prime}(a)\coloneqq f^{\partial}(a)\cdot g(-a)$ is the dual of the unary discriminator. We have shown in (Düntsch et al., 2023) that for binary modalities such equivalence does not hold any more, and the weaker condition
@@
Clearly, every wMIA satisfies ($\boldsymbol{\mathsf{di}}$). On the other hand, Table 1 of (Düntsch et al., 2023) shows that (wMIA) is not equivalent to ($\boldsymbol{\mathsf{di}}$).
@@
Secondly, let $a\neq\mathbf{1}$, i.e., $-a\cdot\mathbf{1}\neq\mathbf{0}$. Then by ($\boldsymbol{\mathsf{di}}$) we have $g(-a,\mathbf{1})\leq f(-a,\mathbf{1})$, and thus
@@
$u$$\langle\mathbf{0},\mathbf{0}\rangle$$\langle\mathbf{0},\mathbf{1}\rangle$$\langle\mathbf{1},\mathbf{1}\rangle$$\langle\mathbf{1},\mathbf{0}\rangle$$\mathbf{1}$$\mathbf{0}$$\Delta$$\mathbf{dMIA}$
Figure 3. Assume that $\langle A,\mathord{\leq_{\ell}}\rangle$ is a linearization (i.e., $\leq_{\ell}$ is an extension of $\leq$ which is a linear order compatible with $\leq$) of a dMIA $A$, and take its binary product. Then, according to Lemmas 4.2 and 4.3, the binary operator $u$ always sends three yellow pairs to $\mathbf{1}$, and the violet ones to $\mathbf{0}$.
@@
If $a\neq\mathbf{1}$ and $b\neq\mathbf{1}$, then $-a\neq\mathbf{0}$ and $-b\neq\mathbf{0}$, and by (wMIA), $g(-a,-b)\leq f(-a,-b)$. Therefore,
@@
$u$$\langle\mathbf{0},\mathbf{0}\rangle$$\langle\mathbf{0},\mathbf{1}\rangle$$\langle\mathbf{1},\mathbf{1}\rangle$$\langle\mathbf{1},\mathbf{0}\rangle$$\mathbf{1}$$\mathbf{0}$$\Delta$$\mathbf{wMIA}$
Figure 4. This time let $\langle A,\mathord{\leq_{\ell}}\rangle$ be a linearization of a wMIA $A$, and consider its binary product. Then, according to Lemma 4.4, the binary operator $u$ sends the three yellow pairs to $\mathbf{1}$, and all the remaining ones to $\mathbf{0}$. The less saturated fragment of the rectangle indicates the area of possible differences between the behavior of $u$ for dMIAs and wMIAs.
@@
Indeed, if $x\neq\mathbf{1}$, then $-x\neq\mathbf{0}$ and by ($\boldsymbol{\mathsf{di}}$) we have $g(-x,\mathbf{1})\leq f(-x,\mathbf{1})$, and so: $u(x,\mathbf{0})=g(-x,\mathbf{1})\cdot-f(-x,\mathbf{1})=\mathbf{0}$, as required.
To show that wMIA holds, assume that $a\neq\mathbf{0}$ and $b\neq\mathbf{0}$, i.e., $-a\neq\mathbf{1}$ and $-b\neq\mathbf{1}$. By the above we have $u(-a,\mathbf{0})=\mathbf{0}$ and $u(-b,\mathbf{0})=\mathbf{0}$. By (4.5) and (4.4) (which is a consequence of ($\boldsymbol{\mathsf{di}}$) by Lemma 4.3) $u(-a,-b)=\mathbf{0}+\mathbf{0}=\mathbf{0}$. Thus, $-f(a,b)\cdot g(a,b)=\mathbf{0}$, i.e., $g(a,b)\leq f(a,b)$. Therefore, $\mathfrak{A}$ is a weak MIA.
@@
Using again the algebra in Figure 5 we can see in light of Proposition 4.5 that (4.5) does not entail ($\boldsymbol{\mathsf{di}}$). Also, the algebra from Table 1 of (Düntsch et al., 2023) shows that ($\boldsymbol{\mathsf{di}}$) is strictly weaker than (wMIA), and thus ($\boldsymbol{\mathsf{di}}$) cannot entail (4.5) either. Therefore, we obtain
@@
Indeed, $-{u^{1}}(-a)=-u(-a,\mathbf{0})=u^{\partial}(a,\mathbf{1})={u^{1}}^{\partial}(a)$.
@@
Then, $d(a)=-u(-a,-a)=f(a,a)+-g(a,a)$ is the unary discriminator. From Lemma 4.2, we obtain ($\boldsymbol{\mathsf{di}}$) holds in $\mathfrak{A}$. If in addition $\mathfrak{A}$ satisfies (4.5), then by Proposition 4.5, $\mathfrak{A}\in\mathbf{wMIA}$. ∎
@@
$u(a,\mathbf{0})\leq a$ $u(a,\mathbf{0})\leq u(u(a,\mathbf{0}),\mathbf{0})$ $a\leq u(-u(-a,\mathbf{0}),\mathbf{0})$ $u(a,a)=u(a,\mathbf{0})=u(\mathbf{0},a)$ $u(a,b)=u(a,\mathbf{0})+u(\mathbf{0},b)$$\operatorname{\mathbf{V}}(\mathbf{dMIA})$(4.5)$\operatorname{\mathbf{V}}(\mathbf{wMIA})$
Finally in this section we will show an analogon to (Düntsch et al., 2017, Theorem 8.5). Let $T(\mathsf{Var})$ be the term algebra over the language of $K^{\#}$ with the set $\mathsf{Var}$ of variables. Then, each formula $\varphi(p_{1},\ldots,p_{n})$ of $K^{\#}$ can be regarded as an element of $T(\mathsf{Var})$ via the natural translation $\tau$. If $\mathfrak{A}$ is an algebra with the signature of $K^{\#}$, we say that $\varphi(p_{1},\ldots,p_{n})$ is valid in $\mathfrak{A}$—written as $\mathfrak{A}\models\varphi(p_{1},\ldots,p_{n})$—if $\tau(\varphi(v(p_{1}),\ldots,v(p_{n})))=\mathbf{1}$ for all valuations $v:\mathsf{Var}\to A$.
@@
By a 3-frame, we mean any Kripke frame $\mathfrak{F}\coloneqq\langle W,R\rangle$ in which $R$ is a ternary relation on $W$. In (Düntsch et al., 2023) we asked the following question: is there for each betweenness algebra $\mathfrak{A}$ a 3-frame $\mathfrak{F}$ such that $\mathfrak{A}$ is embeddable into $\operatorname{{\mathsf{Cm}}}(\mathfrak{F})$? Below, we answer this question positively by showing a stronger statement: any weak mixed algebra $\mathfrak{A}$ can be embedded into the complex algebra of a 3-frame, which is the special frame of the canonical frame of $\mathfrak{A}$. To this end we follow the construction developed in Section 3.3
@@
Figure 7. An application of the construction of the special frame to the canonical frame $\operatorname{{\mathsf{Cm}}}(\mathfrak{A})=\langle\operatorname{Ult}(A),Q_{f},S_{g}\rangle$ of the wMIA $\mathfrak{A}$. $A\coloneqq\{\langle\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}\rangle,\langle\mathscr{U}_{1}^{\prime},\mathscr{U}_{2}^{\prime},\mathscr{U}_{3}^{\prime}\rangle\}$, where $\langle\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}\rangle\in Q_{f}\setminus S_{g}$.

### v2

@@ Abstract.
In previous work [6] we introduced and examined the class of betweenness algebras. In the current paper we study a larger class of algebras with binary operators of possibility and sufficiency, the weak mixed algebras. Furthermore, we develop a system of logic with two binary modalities, sound and complete with respect to the class of frames closely related to the aforementioned algebras, and we prove an embedding theorem which solves an open problem from [6].
@@ 1. Introduction
Possibility-sufficiency algebras (PS-algebras), as Boolean algebras expanded with unary operators of possibility and sufficiency, were first studied in [3]. The latter operator is an algebraic counterpart of the window modality, whose provenance can be traced back to [10, 7, 8], and also [14] in the framework of deontic logic. The theory of PS-algebras was refined and developed further in [4] and [5], with particular emphasis on two subclasses of PS-algebras: mixed and weak mixed algebras, which force a relationship between the—otherwise independent—operators of possibility and sufficiency.
@@
Building on the aforementioned works, we initiated in [6] a theory of betweenness algebras (b-algebras), which were thought of as an approach to study most general algebraic and logical properties of the ternary relation of betweenness well-known from ordered geometry. Being ternary, betweenness calls for binary operators, and for this reason, our PS-algebras are Boolean algebras with binary operators of possibility and sufficiency. To guarantee the interplay between the two operators, we focused on weak mixed algebras, which we enriched with equalities expressing the basic and well-motivated properties of betweenness. The approach turned out to be fruitful and allowed for the development of both representation and canonical extensions of algebras arising from betweenness relations. For these, the canonical frames of b-algebras are triples of the form $\langle W,R,S\rangle$, where $S$ and $R$ are ternary relations on $W$ obtained in the standard way from sufficiency and possibility operators, respectively, with $S\subseteq R$. In a certain sense, $R$ and $S$ together simulate the betweenness relation. We showed that this cannot be avoided, in the sense that, in general, the canonical frame of a b-algebra cannot have just one ternary relation which is a betweenness. This was a negative result, as the starting point of our research were precisely frames $\langle W,B\rangle$, where $B$ is a betweenness, and the properties of the complex algebras of such frames gave rise to the axiomatization presented in the paper. However, the negative finding did not exclude existence of a more general ternary relation that could allow for a representation theorem for b-algebras, and finding such a relation was formulated as problem (1) in [6].
@@
With the problem in focus, we concentrate below on a general study of frames $\langle W,R,S\rangle$ with $S\subseteq R$ (wMIA frames). We develop a logic $K^{\#}$ with binary operators of possibility and sufficiency (window modality) that is sound and complete with respect to the class of wMIA frames. To this end, we adapt for our needs the copying method from [13, 7], which allows for the transformation of wMIA frames into frames with a single ternary relation. Further, we study algebraic models of $K^{\#}$, i.e., PS-algebras determined by an equational theory $\Sigma$. We prove that the variety $\operatorname{\mathbf{V}}(\Sigma)$ is precisely the variety generated by the class of weak mixed algebras (which does not form a variety itself).
@@ 2.1. Boolean algebras
is a Boolean congruence relation. Conversely, if $\theta$ is a Boolean congruence on $A$, then $F_{\theta}\coloneqq\{a\in A:a\mathrel{\theta}\mathbf{1}\}$ is a filter on $A$. Furthermore, $\theta_{F_{\theta}}=\theta$ and $F_{\theta_{F}}=F$.
@@ 2.2. PS-algebras
We prepare the description of the congruences on a PS-algebra $\mathfrak{A}$ with a few known facts. The Boolean congruences on $A$ were described earlier in Lemma 2.1. Congruences with respect to a binary necessity operator were characterized in Venema [15].
Lemma 2.2 (15, Theorem 29).
@@
To apply these results, we first define for a sufficiency operator $g:A\times A\to A$ a mapping $g^{*}:A\times A\to A$ by $g^{*}(a,b)=g(-a,-b)$. The operator $g^{*}$ is a binary necessity operator, and a Boolean congruence preserves $g$ if and only if it preserves $g^{*}$, since $g$ and $g^{*}$ are mutually term definable. We now define an auxiliary mapping $u\colon A\times A\to A$ [^1] The name refers to the universal modality which is related to the unary discriminator, see e.g. [7, 9].
@@ 2.3. Complex algebras and canonical frames
The canonical frame of $\mathfrak{A}$ is the structure $\operatorname{{\mathsf{Cf}}}(\mathfrak{A})\coloneqq\langle\operatorname{Ult}A,R_{f},S_{g}\rangle$, where $R_{f},S_{g}$ are ternary relations on $\operatorname{Ult}A$ defined by
@@
($\supseteq$) Let $\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}\in\operatorname{Ult}A$, and $a\in\mathscr{U}_{1},b\in\mathscr{U}_{3}$. If $R_{f}(\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3})$, then, $f(a,b)\in\mathscr{U}_{2}$, and it follows that $h(a,b)=f(a,b)+-g(a,b)\in\mathscr{U}_{2}$. If $\langle\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}\rangle$ is not an element of $S_{g}$, then $g[\mathscr{U}_{1}\times\mathscr{U}_{3}]\cap\mathscr{U}_{2}=\emptyset$, hence, $g(a,b)\not\in\mathscr{U}_{2}$ which implies $g(a,b)\in\mathscr{U}_{2}$. It follows that $h(a,b)=f(a,b)+-g(a,b)\in\mathscr{U}_{2}$. ∎
@@ 3. The binary logic $K^{\#}$
$U]](\varphi,\psi)\coloneqq\Box(\varphi,\psi)\wedge\boxbar(\neg\varphi,\neg\psi).$
@@
[…]
$\vdash\varphi\rightarrow\psi$
$\vdash\Diamond(\varphi,\chi)\rightarrow\Diamond(\psi,\chi)$
$\vdash\varphi\rightarrow\psi$
$\vdash\Diamond(\chi,\varphi)\rightarrow\Diamond(\chi,\psi)$
[…]
@@ 3.1. Relational semantics
[…]
@@
Notation not explained here can be found in [1, Section 5], [6], and [2], in particular, II §10.
@@
and—thanks to these—with $\langle U\rangle\!\rangle$ and $[U]]$ we can associate global modalities, existential and universal, respectively
@@ 3.2. The canonical frame for $\boldsymbol{K^{\#}}$
$U]](\langle U\rangle\!\rangle(\langle U\rangle\!\rangle(\psi,\top),\top),\bot)\in y$
@@
$U]](\langle U\rangle\!\rangle(\psi,\top),\bot)\in y.$
By the same argument $U]](\langle U\rangle\!\rangle(\varphi,\top),\bot)\in y$. Since
$U]](\langle U\rangle\!\rangle(\varphi,\top),\bot)\land[U]](\langle U\rangle\!\rangle(\psi,\top),\bot)\Leftrightarrow[U]](\langle U\rangle\!\rangle(\varphi,\top)\land\langle U\rangle\!\rangle(\psi,\top),\bot),$
@@
$U]](\langle U\rangle\!\rangle(\varphi,\psi),\bot)\in y,\quad\text{i.e.,}\quad\neg\langle U\rangle\!\rangle(\neg\langle U\rangle\!\rangle(\varphi,\psi),\top)\in y.$
@@ 3.3. Special frames and models
Our next aim is to construct—given $\langle W,R,S,v\rangle$—a special model $\langle\underline{W},\underline{R},\underline{S},\underline{v}\rangle$, that is modally equivalent to $\langle W,R,S,v\rangle$. The idea of the construction of $\underline{R}$ and $\underline{S}$ is to “separate” triples $\langle x,y,z\rangle$ which are in the intersection of $R$ and S-S, i.e. to remove those triples which prevent $R=S$ (see Figure 1).
@@
(3.11) $P\subsetneq m(P)\quad$ ${}\text{ and }\quad P^{\prime}\subsetneq m(P^{\prime}),$
@@
$(\supseteq)$ If $\tau\in m(W^{3}\setminus P)$, then there is $\langle x,y,z\rangle\notin P$ such that $\tau\in c(x,y,z)$. If there were $\langle r,s,t\rangle\in m(P)$ such that $\tau\in c(j(r),j(s),j(t))$, then $P$ and $W^{3}\setminus P$ would have a non-empty intersection, which is impossible.
@@
The mixture $m(\mathfrak{F})$ of any wMIA frame $\mathfrak{F}$ can be turned into a special frame via the following construction of $\underline{R}$ and $\underline{S}$ (below, R-R and S-S are the complements of the two relations in $W^{3}$):
@@
Figure 2. $m(R)\setminus m(S)$ is equal to the union of $m(R)\setminus m(S)]\setminus(R\cup R^{\prime})$ and $(R\cup R^{\prime})\setminus(S\cup S^{\prime})$.
@@
$(v(\varphi)\cup v^{\prime}(\varphi))\times\{s\}\times(v(\psi)\cup v^{\prime}(\psi))]\cap\underline{R}\neq\emptyset.$
@@ 4. The class $\operatorname{\mathbf{V}}(\mathbf{wMIA})$
In the case of unary modalities investigated in [5] a unary PS-algebra $\langle A,f,g\rangle$ is a weak MIA if and only if the mapping defined by $u^{\prime}(a)\coloneqq f^{\partial}(a)\cdot g(-a)$ is the dual of the unary discriminator. We have shown in [6] that for binary modalities such equivalence does not hold any more, and the weaker condition
@@
Clearly, every wMIA satisfies ($\boldsymbol{\mathsf{di}}$). On the other hand, Table 1 from [6] shows that (wMIA) is not equivalent to ($\boldsymbol{\mathsf{di}}$).
@@
Secondly, let $a\neq\mathbf{1}$, i.e., $a\cdot\mathbf{1}\neq\mathbf{0}$. Then by ($\boldsymbol{\mathsf{di}}$) we have $g(-a,\mathbf{1})\leq f(-a,\mathbf{1})$, and thus
@@
[…]
Figure 3. Assume that $\langle A,\mathord{\leq_{\ell}}\rangle$ is a linearization (i.e., $\leq_{\ell}$ is an extension of $\leq$ which is a linear order compatible with $\leq$) of a dMIA AA, and take its binary product. Then, according to lemmas 4.2 and 4.3, the binary operator $u$ always sends three black pairs to $\mathbf{1}$, and the gray ones to $\mathbf{0}$.
@@
If $a\neq\mathbf{1}$ and $b\neq\mathbf{1}$, then $a\neq\mathbf{0}$ and $b\neq\mathbf{0}$, and by (wMIA), $g(-a,-b)\leq f(-a,-b)$. Therefore,
@@
[…]
Figure 4. This time let $\langle A,\mathord{\leq_{\ell}}\rangle$ be a linearization of a wMIA AA, and consider its binary product. Then, according to Lemma 4.4, the binary operator $u$ sends the three black pairs to $\mathbf{1}$, and all the remaining ones to $\mathbf{0}$. The less saturated fragment of the rectangle indicates the area of possible differences between the behavior of $u$ for dMIAs and wMIAs.
@@
Indeed, if $x\neq\mathbf{1}$, then $x\neq\mathbf{0}$ and by ($\boldsymbol{\mathsf{di}}$) we have $g(-x,\mathbf{1})\leq f(-x,\mathbf{1})$, and so: $u(x,\mathbf{0})=g(-x,\mathbf{1})\cdot-f(-x,\mathbf{1})=\mathbf{0}$, as required.
To show that wMIA holds, assume that $a\neq\mathbf{0}$ and $b\neq\mathbf{0}$, i.e., $a\neq\mathbf{1}$ and $b\neq\mathbf{1}$. By the above we have $u(-a,\mathbf{0})=\mathbf{0}$ and $u(-b,\mathbf{0})=\mathbf{0}$. By (4.5) and (4.4) (which is a consequence of ($\boldsymbol{\mathsf{di}}$) by Lemma 4.3) $u(-a,-b)=\mathbf{0}+\mathbf{0}=\mathbf{0}$. Thus, $f(a,b)\cdot g(a,b)=\mathbf{0}$, i.e., $g(a,b)\leq f(a,b)$. Therefore, $\mathfrak{A}$ is a weak MIA.
@@
Using again the algebra in Figure 5 we can see in light of Proposition 4.5 that (4.5) does not entail ($\boldsymbol{\mathsf{di}}$). Also, the algebra from Table 1 of [6] shows that ($\boldsymbol{\mathsf{di}}$) is strictly weaker than (wMIA), and thus ($\boldsymbol{\mathsf{di}}$) cannot entail (4.5) either. Therefore, we obtain
@@
Indeed, ${u^{1}}(-a)=-u(-a,\mathbf{0})=u^{\partial}(a,\mathbf{1})={u^{1}}^{\partial}(a)$.
@@
Then, d(a)=u(a,a)=f(a,a)+g(a,a)d(a)=-u(-a,-a)=f(a,a)+-g(a,a) is the unary discriminator. From Lemma 4.2, we obtain ($\boldsymbol{\mathsf{di}}$) holds in $\mathfrak{A}$. If in addition $\mathfrak{A}$ satisfies (4.5), then by Proposition 4.5, $\mathfrak{A}\in\mathbf{wMIA}$. ∎
@@
[…]
Finally in this section we will show an analogon to [5, Theorem 8.5]. Let $T(\mathsf{Var})$ be the term algebra over the language of $K^{\#}$ with the set $\mathsf{Var}$ of variables. Then, each formula $\varphi(p_{1},\ldots,p_{n})$ of $K^{\#}$ can be regarded as an element of $T(\mathsf{Var})$ via the natural translation $\tau$. If $\mathfrak{A}$ is an algebra with the signature of $K^{\#}$, we say that $\varphi(p_{1},\ldots,p_{n})$ is valid in $\mathfrak{A}$—written as $\mathfrak{A}\models\varphi(p_{1},\ldots,p_{n})$—if $\tau(\varphi(v(p_{1}),\ldots,v(p_{n})))=\mathbf{1}$ for all valuations $v:\mathsf{Var}\to A$.
@@
By a 3-frame, we mean any Kripke frame $\mathfrak{F}\coloneqq\langle W,R\rangle$ in which $R$ is a ternary relation on $W$. In [6] we asked the following question: is there for each betweenness algebra $\mathfrak{A}$ a 3-frame $\mathfrak{F}$ such that $\mathfrak{A}$ is embeddable into $\operatorname{{\mathsf{Cm}}}(\mathfrak{F})$? Below, we answer this question positively by showing a stronger statement: any weak mixed algebra $\mathfrak{A}$ can be embedded into the complex algebra of a 3-frame, which is the special frame of the canonical frame of $\mathfrak{A}$. To this end we follow the construction developed in Section 3.3
@@
Figure 7. An application of the construction of the special frame to the canonical frame $\operatorname{{\mathsf{Cf}}}^{ps}(\mathfrak{A})=\langle\operatorname{Ult}A,Q_{f},S_{g}\rangle$ of the wMIA $\mathfrak{A}$. $M\coloneqq\{\langle\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}\rangle,\langle\mathscr{U}_{1}^{\prime},\mathscr{U}_{2}^{\prime},\mathscr{U}_{3}^{\prime}\rangle\}$, where $\langle\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}\rangle\in Q_{f}\setminus S_{g}$.